Grade 11Mechanics

at t=0 a particle at origin has a velocity 15m/s at 37° above the horizontal x axis. at t =5s it is at x=20m and y=35m and its velocity is 30 m/s at 53° above the horizontal. Find its average acceleration

garvit
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
Time of Flight for case first:
T1 = 2*u*sinthetha/g = > 2*15*sin37/g = 18/g
Time of Flight for case second:
T2 = 2vsinthetha/g = > 2*30sin53/g = 48/g
 
average acceleration,
= (v – u)/(T2 – T1)
= (30 – 15)/(48/g –  18/g)
= 15/30g
=1/2g
= 1/2*9.8 m/s^2